OpenStudy (anonymous):

6 miles away from her and 4 miles above from the ground. how many degrees horizontal must she look up to see the plane?

OpenStudy (anonymous):

Pythagoras' theorem: \[c^2 = a^2 + b^ 2\] so \[c^2 = 4^2 + 6^2\] \[c^2 = 16 + 36 = \] \[c = \sqrt{52} \approx 7.21\] \[\sin^{-1} \alpha = \frac{ 4}{ 7.21 } \rightarrow \alpha = 33.69^{o}\]